OpenStudy (anonymous):

Estimate the tip on a $19.65 check using a tip rate of 15%

OpenStudy (anonymous):

Dude I just answered this.

OpenStudy (anonymous):

What is 15% of $19.65? Usually when you see the word "of" it indicates multiplication. You do: $19.65 x 0.15 = 2.9475 <-- This is the 15% out of the total $19.65 and is how much you paid for tips.

jimthompson5910 (jim_thompson5910):

You can round 19.65 up to 20 and then multiply 15 by 20 to get 300, but you have to move the decimal back 3 spots to get 3 So the tip is roughly $3

jimthompson5910 (jim_thompson5910):

of course, it's a bit smaller than that, but that's just an estimate

OpenStudy (anonymous):

You multiply 0.15 , not 15 because we are dealing with percentage. 100% = 100/100= 1.00 15% = 15/100 = 0.15

jimthompson5910 (jim_thompson5910):

I know, I'm ignoring the decimals at first, then moving decimal points

OpenStudy (anonymous):

Or $2.95 when dealing with money

OpenStudy (anonymous):

thNks